Question 1:

payment frequency = weekly

P/Y=52, C/Y=2, I/Y=3.75, PV=383427, FV=0.

Step 1: Calculate PMT.

Step 2: Calculate N(number of periods)

Step 3: Calculate AMORT =N / (P/Y)
Answer #1

PMT = PMT (Rate, Nper, PV, FV) = PMT (3.75% / 52, 2 x 52, -383427, 0) = $  3,828.11

N = C/Y x P/Y = 2 x 52 = 104

AMORT = N / (P/Y) = 104 / 52 = 2